# AttnLRP: Attention-Aware Layer-wise Relevance Propagation for Transformers

### v1

@@ A.1.5 AtMan
$\textbf{H}=\textbf{Q}\cdot\textbf{K}^{T}$
@@
where $\textbf{H}\in\mathbb{R}^{b\times q\times k}$, and $\textbf{1}\in[1]^{b\times q\times k}$ a matrix containing only 1. $\textbf{p}^{i}$ denotes a matrix $\in\mathbb{R}^{b\times q\times k}$ with
@@ A.2.1 Conservation & Numerical Stability of Bias Terms
The relevance of a specific neuron or function, denoted as $R_{j}$, is computed by summing the contributions of the input variables, represented by $\textbf{J}_{ji}\ x_{i}$, and adding the contribution of the bias term, represented by $\tilde{b_{j}}$.
$R_{j}=\left(\sum_{i}\textbf{J}_{ji}\ x_{i}+\tilde{b_{j}}\right)\frac{R_{j}}{f_{j}(\textbf{x})}$
The relevance of the bias term itself, denoted as $R_{b}$, is calculated as:
$R_{b}=\tilde{b_{j}}\frac{R_{j}}{f_{j}(\textbf{x})}$
If we want to compute the relevance of the input variables $R_{i}$ while ensuring strict adherence to the conservation property (3), we must exclude the bias term, so that it does not absorb part of the relevance $R_{b}$. There are three options: We can either completely omit the bias term, distribute its relevance value uniformly across the input variables, or apply the identity rule.
Omitting the bias term: In this case, the relevance propagation equation is:
$R_{i}=\sum_{j}R_{i\leftarrow j}=\sum_{j}\textbf{J}_{ji}\ x_{i}\frac{R_{j}}{\sum_{i}\textbf{J}_{ji}\ x_{i}+\varepsilon}$
Here, we no longer divide by the original function $f_{j}(\textbf{x})$. However, it is important to ensure that no sign flips occur, as $\sum_{i}\textbf{J}_{ji}\ x_{i}$ might have a different sign than $f_{j}(\textbf{x})$.
Distributing the bias term: Alternatively, we can distribute the relevance value of the bias term uniformly across the input variables. The relevance propagation equation in this case is:
$R_{i}=\sum_{j}R_{i\leftarrow j}=\sum_{j}\left(\textbf{J}_{ji}\ x_{i}+\frac{\tilde{b_{j}}}{N}\right)\frac{R_{j}}{f_{j}(\textbf{x})+\varepsilon}$
Here, $N$ represents the number of input variables and $R_{b}=0$. However, this approach may lead to numerical instabilities if the function $f_{j}(\textbf{x})=0$ with $x_{i}=0$. Then, the relevance message $R_{i\leftarrow j}$ is:
$R_{i\leftarrow j}=\frac{\tilde{b_{j}}}{N}\frac{R_{j}}{\varepsilon}$
Since $\varepsilon$ is very small, this term explodes and causes numerical instabilities. Thus, functions that fulfill $f(0)=0$ should preferably linearized such that the bias term does not cause numerical instabilities.
Applying the identity rule: We can apply the identity rule as follows:
@@
However, we may encounter numerical instabilities, but these effects will only become visible in the next sequential propagation rule, not at this layer yet. For example in the softmax function, we may encounter a situation where a $x_{i}^{l-1}=0$ but $f_{i}^{l}(\textbf{x})>0$. If a non-zero relevance value from layer $l$ is assigned to $f_{i}^{l}(\textbf{x})$, then its relevance value is propagated through the identity rule to $x_{i}^{l-1}$. Assuming we apply the $\varepsilon$-rule after the identity rule, the relevance is given by:
Note, that we divide by $x_{i}^{l-1}=f_{i}^{l-1}=0$, which results in numerical instabilities. Further note, that these instabilities would not occur if $x_{k}^{l-2}=0$ or $R_{i}^{l-1}=0$, which is the case for a function fulfilling $f^{l}_{i}(0)=0$.
In summary, omitting the bias term completely, distributing its relevance value uniformly across the input variables or applying the identity rule are possible approaches, but they have their considerations and potential challenges. Regarding the softmax non-linearity, Voita et al., (2021) distributes the bias term equally on all input variables, while Ding et al., (2017); Chefer et al., 2021b apply the element-wise identity rule (9). Both variants can lead in this case to severe numerical instabilities.
@@ A.2.2 Highlighting the difference between various LRP methods
In Table 3, we illustrate the different strategies employed for LRP in the past.
As discussed in Appendix A.2.1, distributing the bias term uniformly or applying the identity rule on the softmax function can lead to numerical instabilities that we also observed in our experiments.
Applying the $\varepsilon$-rule on bi-linear matrix multiplication does violate the conservation property (3) as proved in Lemma 3 of Chefer et al., 2021b . As a heuristic, Chefer et al., 2021b applies a normalization step by dividing both arguments by the summation of its absolute values.
Ali et al., (2022) regards the softmax output as constant and does not propagate relevance through it. This way, the matrix multiplication is not bi-linear anymore, but becomes linear, resulting in the attribution of only the value path. Consequently, the query and key matrices can no longer be attributed, which reduces the faithfulness and makes latent explanations in query and key matrices infeasible.
@@
In Figure 5, we illustrate different attribution maps for all four options to handle the softmax function. The given section is from the Wikipedia article on Mount Everest. The model is expected to provide the next answer token for the question ‘How high did they climb in 1922? According to the text, the 1922 expedition reached 8,’. For the correctly predicted token 3 is the attribution computed.
@@
While the relevance values for AttnLRP or CP-LRP are between $[-4,4]$, distributing the bias uniformely on the input variables or applying the identity rule leads to an explosion of the relevances between $[-10^{15},10^{15}]$. As a consequence, the heatmaps resemble random noise. AttnLRP highlights the correct token the strongest, while CP-LRP focuses strongly on the start-of-sequence <s> token and exhibits more background noise e.g. irrelevant tokens such as ‘Context’, ‘attracts’, ‘Everest’ are highlighted, while AttnLRP does not highlight them or assigns negative relevance. In Appendix B.5, we compare also other baseline methods. Note, that the model attends to numerous tokens within the text which enables it to derive conclusions. Consequently, an attribution that reflects the model behavior will highlight more than just the single accurate answer token. The faithfulness experiments in Table 1 demonstrate, the AttnLRP captures the model reasoning most accurately.
@@
Table 3: Conceptual differences between various-LRP methods and their implications.
@@
Ding et al., (2017) Identity rule $\varepsilon$-LRP not available
@@
Voita et al., (2021) Taylor Decomposition at x $\varepsilon$-LRP Taylor Decomposition at x
(distributes the bias uniformly) (distributes the bias uniformly)
@@
Chefer et al., 2021b Identity rule $\varepsilon$-LRP Identity rule
$\Rightarrow$ unstable (Appendix A.2.1) (with absolute normalization)
Ali et al., (2022) Regarded as constant $\varepsilon$-LRP Identity rule
$\Rightarrow$ No attribution (ensures conservation)
inside attention module
@@
(ensures conservation)
@@ A.2.3 Tackling Noise in Vision Transformers
by only taking into account positive contributions $z_{j}^{+}=\sum_{i}(w_{ij}x_{i})^{+}$ with $(\cdot)^{+}=\max(0,\cdot)$. This is the case of softmax layers (of the attention layer) where only positive values are dealt with.
Remarkably, our observations reveal that attributions in LLMs demonstrate high sparsity and lack visible noise, while ViTs are susceptible to gradient shattering. We hypothesize that the discrete nature of the text domain may affect robustness Mao et al., (2021). Therefore, we only apply the $\gamma$-rule in ViTs in the convolutional and linear FFN layers outside the attention module. To further increase the faithfulness, the $\gamma$-rule can be also applied on softmax layers. Since the output of the softmax is always greater than zero, we apply the simplified $z^{+}$-rule (special case of $\gamma$-rule).

### v2

@@ A.1.5 AtMan
$\textbf{H}=\textbf{Q}\cdot\textbf{K}^{\top}$
@@
where $\textbf{H}\in\mathbb{R}^{b\times s_{q}\times s_{k}}$, and $\textbf{1}\in[1]^{b\times s_{q}\times s_{k}}$ a matrix containing only 1. $\textbf{p}^{i}$ denotes a matrix $\in\mathbb{R}^{b\times s_{q}\times s_{k}}$ with
@@
A.1.6 KernelSHAP
LIME computes attributions by fitting an additive surrogate model Ribeiro et al., (2016). KernelSHAP Lundberg and Lee, (2017) is a special case of LIME, that sets the loss function, weighting kernel and regularization terms of LIME such that LIME recovers Shapley values. Hence, KernelSHAP allows theoretically to obtain Shapley Values more efficiently than directly computing Shapley Values.
To apply KernelSHAP in the vision domain, we divide the input image into $N$ super-pixels using the Simple Linear Iterative Clustering (SLIC) algorithm Achanta et al., (2012). We use captum Kokhlikyan et al., (2020) with its default settings to compute the attributions i.e. number of samples per attributions set to 2000 and baseline value set to 0.5. A baseline value of 0 resulted in lower faithfulness. For SLIC, we set $N=100$ with compactness set to 10.
@@ A.2.1 Conservation & Numerical Stability of Bias Terms
The total relevance $R_{j}$ of a layer output, i.e. linearized function $f_{j}(\textbf{x})=\sum_{i}\textbf{J}_{ji}\ x_{i}+\tilde{b}_{j}$, is computed by summing the contributions of the input variables $R_{i\leftarrow j}$, represented by $\textbf{J}_{ji}\ x_{i}$, and adding the contribution of the bias term $R_{b\leftarrow j}$, represented by $\tilde{b}_{j}$.
$R_{j}=\sum_{i}R_{i\leftarrow j}+R_{b\leftarrow j},$
The relevance of the input variables is solely determined by the input variables themselves
$R_{i}=\sum_{j}R_{i\leftarrow j}=\sum_{j}\textbf{J}_{ji}\ x_{i}\frac{R_{j}}{f_{j}(\textbf{x})},$
while the relevance of the bias term itself is calculated as
$R_{b\leftarrow j}=\tilde{b}_{j}\frac{R_{j}}{f_{j}(\textbf{x})}.$
If we want to compute the relevance of the input variables $R_{i}$ while ensuring strict adherence to the conservation property (3), we must exclude the bias term, so that it does not absorb part of the relevance. In the literature, we find two common practices: Either distributing the bias term uniformely on the input variables Binder et al., (2016); Voita et al., (2021) or applying the identity rule Ding et al., (2017); Chefer et al., 2021b . Both approaches can lead to severe numerical instabilities in specific cases that are challenging to identify. Therefore, we will dedicate some time to explain the issue in greater detail.
Remark A.2.1 Enforcing strict conservation (3) on a function, where
$\exists i,j\in\mathbb{N}:x_{i}=0\land f_{j}(\textbf{x})\neq 0,$
by distributing the relevance of the bias term of its linearization uniformly on the input variables or applying the identity rule with $i=j$ may lead to numerical instabilities.
Distributing the bias term: We can distribute the relevance value of the bias term uniformly across the input variables by assuming that the bias term is part of the input variables:
$R_{j}^{l}=\sum_{i}\tilde{R}_{i\leftarrow j}^{(l-1,l)}=\sum_{i}^{N}\left(R_{i\leftarrow j}^{(l-1,l)}+\frac{R_{b\leftarrow j}^{(l-1,l)}}{N}\right),$
where $N$ represents the number of input variables. Hence,
$R_{i}^{l}=\sum_{j}\tilde{R}_{i\leftarrow j}^{(l-1,l)}=\sum_{j}\left(\textbf{J}_{ji}\ x_{i}+\frac{\tilde{b_{j}}}{N}\right)\frac{R_{j}^{l}}{f_{j}(\textbf{x})}.$
However, we may encounter numerical instabilities, but these effects will only become visible in the next sequential relevance propagation at the prior layer, not at this layer yet. For example in the softmax function, we may encounter a situation where $\exists i,j\in\mathbb{N}$ with $x_{i}^{l-1}=0$ but $f_{j}^{l}(\textbf{x})>0$. If a non-zero relevance value from layer $l$ is assigned to $f_{j}^{l}(\textbf{x})$, then its relevance value is propagated to the input variable $x_{i}^{l-1}$ through the relevance message:
$\tilde{R}_{i\leftarrow j}^{(l-1,l)}=\frac{\tilde{b}_{j}}{N}\frac{R_{j}^{l}}{f_{j}^{l}(\textbf{x})+\varepsilon}$
Assuming we apply the $\varepsilon$-rule in succession, the relevance in the prior layer is given by:
Here, we divide by $x_{i}^{l-1}=f_{i}^{l-1}=0$. Since $\varepsilon$ is very small, this term explodes and causes numerical instabilities. Hence, assigning a non-zero relevance value to an input that equals zero leads to numerical instabilities. Note, that these instabilities would not occur if $R_{i}^{l-1}=0$ e.g. $R_{i\leftarrow j}^{(l-1,l)}=0$. Functions where $f(0)=0$ do not encounter this issue, because zero output activations will not receive any relevance in following layers e.g. $R_{j}^{l}=0$ using all rules described in this paper.
Applying the identity rule: Alternatively, we can apply the identity rule as follows:
@@
Here, numerical instabilities might also arise in the subsequent relevance propagation, not at this layer. With the same reasoning as before, the identity rule propagates a non-zero relevance value to an input variable that is zero.
Omitting the bias term: For the sake of completeness, we mention that omitting the bias term entirely is also an option. In this case, the relevance propagation equation is:
$R_{i}^{l-1}=\sum_{j}\textbf{J}_{ji}\ x_{i}\frac{R_{j}^{l}}{\sum_{i}\textbf{J}_{ji}\ x_{i}+\varepsilon}$
Here, we no longer divide by the original function $f_{j}(\textbf{x})$, but by its linearization without the bias term. However, it is important to ensure that no sign flips occur, as $\sum_{i}\textbf{J}_{ji}\ x_{i}$ might have a different sign than $f_{j}(\textbf{x})$.
Remark A.2.2 Enforcing strict conservation (3) by omitting the bias term of a linearization (3.1.1) can lead to sign flips in the relevance scores.
Summary: In summary, applying the identity rule, distributing its relevance value uniformly across the input variables or omitting the bias term completely are possible approaches, but they have their considerations and potential challenges. Regarding the softmax non-linearity, Voita et al., (2021) distributes the bias term equally on all input variables, while Ding et al., (2017); Chefer et al., 2021b apply the element-wise identity rule (9). Both variants can lead to severe numerical instabilities.
@@ A.2.2 Highlighting the difference between various LRP methods
In Table LABEL:table:lrpcomparison, we illustrate the different strategies employed for LRP in the past.
Softmax: Voita et al., (2021) linearizes at x but distributes the bias term equally on all input variables, while Ding et al., (2017); Chefer et al., 2021b apply the element-wise identity rule (9). More specifically, Ding et al., (2017) did not discuss the softmax function explicitly, but they skip all non-linear activation functions. Therefore, we assume that Ding et al., (2017) applies the identity rule also to the softmax function. Both variants enforce a strict notion of the conservation principle (3), but can lead to severe numerical instabilities as discussed in Appendix A.2.1. Ali et al., (2022) regards the attention matrix A in Equation (11) as constant, attributing relevance solely through the value path by stopping the relevance flow through the softmax. Consequently, the query and key matrices can no longer be attributed, which reduces the faithfulness and makes latent explanations in query and key matrices infeasible. Finally, AttnLRP linearizes at x with a bias term that absorbs part of the relevance. The presence of a bias term in AttnLRP is justified because the softmax function yields a value of 1/N even when the input is zero. This is analogous to a bias term and is necessary to account for this behavior. This ensures not only numerically robust attributions, but also improves the faithfulness considerably.
@@
In Figure A.4, we illustrate different attribution maps for all four options to handle the softmax function. The given section is from the Wikipedia article on Mount Everest. The model is expected to provide an answer for the question ‘How high did they climb in 1922?’ and for the correctly predicted next token 3 of the answer ‘According to the text, the 1922 expedition reached 8,’ is the attribution computed by initializing the relevance at the predicted token with its logit value.
@@
While the relevance values for AttnLRP or CP-LRP are between $[-4,4]$, distributing the bias uniformely on the input variables or applying the identity rule leads to an explosion of the relevances between $[-10^{15},10^{15}]$. As a consequence, the heatmaps resemble random noise. AttnLRP highlights the correct token the strongest, while CP-LRP focuses strongly on the start-of-sequence <s> token and exhibits more background noise e.g. irrelevant tokens such as ‘Context’, ‘attracts’, ‘Everest’ are highlighted, while AttnLRP does not highlight them or assigns negative relevance. In Appendix B.7, we compare also other baseline methods. Note, that the model attends to numerous tokens within the text which enables it to derive conclusions. Consequently, an attribution that reflects the model behavior will highlight more than just the single accurate answer token. The faithfulness experiments in Table 1 demonstrate, that AttnLRP captures the model reasoning most accurately.
Matrix Multiplication: Applying the $\varepsilon$-rule (8) on bi-linear matrix multiplication (11) violates the conservation property (3) as proved in Appendix A.3.5. To the best of our knowledge, Ding et al., (2017) applies the standard $\varepsilon$-rule. Voita et al., (2021) utilizes the $z^{+}$-rule (25), that similar to the $\varepsilon$-rule also violates the conservation property (3) in bi-linear matrix multiplication (proof in Appendix A.3.5 is valid for $z^{+}$-rule). While Chefer et al., 2021b also applies the $\varepsilon$-rule, an additional normalization step is performed by dividing both arguments by the summation of its absolute values. This ensures conservation but is not conform with the DTD framework. Ding et al., (2017); Chefer et al., 2021b set the $\varepsilon$ parameter to 00, which may increases numerical instabilities. Hence, we call their LRP variants in Table LABEL:table:lrpcomparison 00-LRP.
Since Ali et al., (2022) regards the softmax output as constant and does not propagate relevance through it, the matrix multiplication is not bi-linear anymore, but becomes linear. Hence, the application of the $\varepsilon$-rule does not violate the conservation principle and attributes only the value path. Ali et al., (2022) sets the $\varepsilon$ parameter to zero, hence we call their LRP variant in Table LABEL:table:lrpcomparison 00-LRP.
Finally, AttnLRP also applies the $\varepsilon$-rule on bi-linear matrix multiplication. In addition, a novel uniform rule (14) derived from the DTD framework is incorporated that ensures conservation and high faithfulness.
Layer Normalization: The works Chefer et al., 2021b ; Ali et al., (2022) and AttnLRP apply the identity rule on normalization functions (18), while using the $\varepsilon$-rule on all linear components of LayerNorm, if applicable. More specifically, Ali et al., (2022) proposes to regard $g(\textbf{x})$ in (18) as a constant, which transforms the normalization operation, and hence the complete LayerNorm layer, into a linear layer, on which the $\varepsilon$-rule is applied. However, this is similar to applying the identity rule (9) on the normalization itself, because it becomes element-wise with a single input and output variable (see Section 3.2.2), while applying the $\varepsilon$-rule on all other linear components of LayerNorm. Voita et al., (2021) linearizes at x and distributes the bias term equally on all input variables, which can lead to numerical instabilities as discussed in Appendix A.2.1.
Vision Transformer: The studies by Ding et al., (2017); Voita et al., (2021); Ali et al., (2022) concentrate on the attribution in natural language processing (NLP) models and do not address vision transformers. Their methodologies, as demonstrated in Table 1 (and Appendix A.2.1), fail when implementing the $\varepsilon$-rule, leading to gradient shattering and low faithfulness. To mitigate noisy attributions, Chefer et al., 2021b suggests employing on-top of LRP an attention rollout Abnar and Zuidema, (2020) procedure which is additionally enhanced via gradient weighting. This yields an approximation of the mean squared relevance value, which diverges from the originally defined notion of “relevance” or “importance” of additive explanatory models. Subsequent empirical observations by Chefer et al., 2021a revealed that an omission of LRP-inspired relevances and a sole reliance on a positive mean-weighting of the attention’s activation with the gradient improved the faithfulness. Though, this approach can only attribute positively, does not consider counteracting evidence, and does not allow to attribute latent neurons outside the attention’s softmax output. AttnLRP, in contrast, adopts the $\gamma$-rule instead of the $\varepsilon$-rule in linear layers, achieving highly faithful attributions without the necessity for a rollout mechanism. However, the $\gamma$ parameter must be tuned to the model and dataset to obtain optimal attributions, as discussed in Appendix A.2.3.
@@
Table A.3: Conceptual differences between various-LRP methods and their implications. “Layer Normalization” refers here only to the normalization function (18) itself and not to the learnable parameters of LayerNorm or RMSNorm.
@@
Ding et al., (2017) Identity rule 00-LRP not available
(bi-linear)
@@
Voita et al., (2021) Taylor decomposition at x $z^{+}$-LRP Taylor decomposition at x
(distributes the bias uniformly) (bi-linear) (distributes the bias uniformly)
@@
Chefer et al., 2021b Identity rule 00-LRP Identity rule
& post-hoc normalization
(bi-linear)
$\Rightarrow$ unstable (Appendix A.2.1) $\Rightarrow$ ensures conservation $\Rightarrow$ ensures conservation
& faithful
Ali et al., (2022) Regarded as constant 00-LRP Identity rule
(linear only)
$\Rightarrow$ stable & no attribution $\Rightarrow$ ensures conservation $\Rightarrow$ ensures conservation
inside attention module & faithful
@@
(bi-linear)
$\Rightarrow$ stable & faithful $\Rightarrow$ ensures conservation $\Rightarrow$ ensures conservation
& faithful & faithful
@@ A.2.3 Tackling Noise in Vision Transformers
by only taking into account positive contributions $z_{j}^{+}=\sum_{i}(w_{ij}x_{i})^{+}$ with $(\cdot)^{+}=\max(0,\cdot)$.
Remarkably, our observations reveal that attributions in LLMs demonstrate high sparsity and lack visible noise, while ViTs are susceptible to gradient shattering. We hypothesize that the discrete nature of the text domain may affect robustness Mao et al., (2021). Therefore, we only apply the $\gamma$-rule in ViTs in the convolutional and linear FFN layers outside the attention module. To further increase the faithfulness, the $\gamma$-rule can be also applied on softmax layers. Since the output of the softmax is always greater than zero, we apply the simplified $z^{+}$-rule (special case of $\gamma$-rule). The $z^{+}$-rule applied on a linearization (3.1.1) for softmax results in:
$R_{i}^{l-1}=\sum_{j}(\textbf{J}_{ji}\ x_{i})^{+}\frac{R_{j}^{l}}{\sum_{k}(\textbf{J}_{jk}\ x_{k})^{+}+\tilde{b}_{j}^{+}}$ (26)
This formula is computationally more expensive to evaluate than the original rule for softmax derived in Proposition 3.1. Should efficiency be a priority, it is recommended to bypass the softmax layer as proposed in CP-LRP, which prevents relevance from passing through the softmax function and reduces gradient shattering caused by this layer. Then, for all other components of the model, AttnLRP rules are recommended, with the application of the $\gamma$-rule to linear layers only. This is especially true, given that the discrepancy in faithfulness between AttnLRP and $\gamma$CP-LRP is minimal for standard vision architectures Dosovitskiy et al., (2021) evaluated in Table 1 and Table B.6, that incorporate only standard attention and FFN layers.
A.2.4 Impact of Temperature Scaling on the Softmax Rule
Temperature scaling controls the entropy within the softmax probability distribution, thereby influencing the predictability of subsequent next token predictions at the classification output. A high temperature value tends to flatten the softmax output distribution (more randomness), whereas a small temperature parameter sharpens the distribution (less randomness). This scaling is done by dividing the input x by the temperature $T\in\mathbb{R}$ prior to applying the softmax function.
$s_{j}(\textbf{x})=\frac{e^{x_{j}/T}}{\sum_{i}e^{x_{i}/T}}$
Recall that the derivative of the softmax function has two cases, which depend on the output $j$ and input $i$ indices:
In the scenario where $s_{j}\approx 1$, the derivative for $i=j$ vanishes. This occurs e.g. for extremely low temperature values or exceptionally high confidence in the model’s classification output. This poses an issue for the Deep Taylor Decomposition Montavon et al., (2017) derived in Section 3.1.1, because DTD decomposes the softmax function by utilizing the gradient (jacobian) term $\textbf{J}_{ji}x_{i}$ for calculating attributions. If the gradient vanishes, the bias term will capture all the relevance, stopping the relevance flow altogether. This effect is also generally described by Shrikumar et al., (2017).
Within the attention mechanism, this limitation is circumvented by multiplying the softmax output with the value path, ensuring that relevance is transmitted via the uniform rule to the value path, akin to CP-LRP (refer to Appendix A.2.2). However, in instances where the softmax function is utilized independently, this becomes problematic as the relevance flow could be distorted.
To see this, consider Proposition 3.1 (13): $R^{l-1}_{i}$ might be zero if $R^{l}_{j}=0\ \forall j\neq i$ with $s_{i}=1$:
$R^{l-1}_{i}=x_{i}(R^{l}_{i}-s_{i}\sum_{j}R^{l}_{j})=x_{i}(R^{l}_{i}-R^{l}_{i})=0$
Therefore, we suggest utilizing an increased temperature scaling value when explaining the softmax classification output to prevent that the softmax saturates i.e. the gradient vanishes. Nonetheless, the attribution of the classification output has not been investigated in this work (the softmax layer is always removed and LRP only applied to the logit outputs). An analysis of these effects remain an interesting topic for future work.